| #1387 |
31 Mar 11:33 Log created ISCOD HLStormtime 31 Mar - 04 Apr
|
ryanque
|
General
|
normal
|
ISR
|
|
| #1388 |
31 Mar 17:00 Antenna Setup
|
walter rojas
|
Antennas
|
normal
|
ISR
|
|
| #1389 |
31 Mar 17:15 Start Experiment
|
ivan manay
|
Start experiment
|
normal
|
ISR
|
|
| #1390 |
31 Mar 17:34 UE Channel C
|
ivan manay
|
General
|
normal
|
ISR
|
|
| #1391 |
31 Mar 17:40 Amplification Stage on Faraday
|
ivan manay
|
Reception
|
normal
|
ISR
|
|
| #1392 |
31 Mar 17:50 Attenuation on Faraday's Channels
|
ivan manay
|
Start experiment
|
normal
|
ISR
|
|
| #1393 |
31 Mar 17:55 Faraday's Plots
|
dsuarez
|
Start experiment
|
normal
|
ISR
|
|
| #1394 |
31 Mar 19:50 Transmitter Power measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1395 |
31 Mar 17:55 IDL program is hung
|
dsuarez
|
Process/Analysis Program
|
normal
|
ISR
|
|
| #1396 |
31 Mar 19:23 Power Amplifier 1 had problems
|
dsuarez
|
Transmitters
|
normal
|
ISR
|
|
| #1397 |
31 Mar 21:07 Spark Gaps adjustment
|
AValdez
|
Spark Gaps
|
normal
|
ISR
|
|
| #1398 |
31 Mar 22:51 Power Amplifier 2 fall
|
AValdez
|
Transmitters
|
normal
|
ISR
|
|
| #1399 |
31 Mar 23:04 Power Amplifier 2 fall
|
AValdez
|
Transmitters
|
normal
|
ISR
|
|
| #1400 |
01 Apr 01:11 Spark Gaps adjustment
|
AValdez
|
Spark Gaps
|
normal
|
ISR
|
|
| #1401 |
01 Apr 05:19 Spark Gaps adjustment
|
AValdez
|
Spark Gaps
|
normal
|
ISR
|
|
| #1402 |
01 Apr 06:20 DP_Faraday Acquisition System stopped
|
AValdez
|
Acquisition system
|
normal
|
ISR
|
|
| #1403 |
31 Mar 20:00 Transmitter Power mesearuments
|
|
Transmitters
|
normal
|
ISR
|
Others
|
| #1404 |
01 Apr 07:13 DP_Faraday Acquisition System stopped
|
AValdez
|
Acquisition system
|
normal
|
ISR
|
|
| #1405 |
01 Apr 08:45 Spark Gaps adjustment
|
I. Manay
|
Spark Gaps
|
normal
|
ISR
|
|
| #1406 |
01 Apr 09:25 Spark gap adjusment
|
I. Manay
|
Spark Gaps
|
normal
|
ISR
|
|
| #1407 |
01 Apr 13:25 Spark Gaps adjustment
|
I. Manay
|
Spark Gaps
|
normal
|
ISR
|
|
| #1408 |
01 Apr 16:20 Spark Gaps adjustment
|
I. Manay
|
Spark Gaps
|
normal
|
ISR
|
|
| #1409 |
01 Apr 19:35 Re-starting the pc of ew_drift experiment
|
I. Manay
|
Acquisition system
|
normal
|
ISR
|
|
| #1410 |
01 Apr 19:50 Transmitter Power measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1411 |
01 Apr 20:41 Calibration ATrs and Trs
|
A. Coronado
|
Antennas
|
normal
|
ISR
|
|
| #1412 |
01 Apr 21:21 Difference in Noise of Ch A and B is more of 2 db
|
P.Condor
|
Reception
|
normal
|
ISR
|
|
| #1413 |
01 Apr 21:25 ChA is attenuated at 1 db
|
P. Condor
|
Reception
|
normal
|
ISR
|
|
| #1414 |
01 Apr 21:35 ChA is attenuated at 1 db
|
P. Condor
|
Reception
|
normal
|
ISR
|
|
| #1415 |
01 Mar 08:00 Transmitter Power mesearuments
|
|
Transmitters
|
normal
|
ISR
|
|
| #1416 |
01 Apr 22:22 Change Tr at ChA
|
P. Condor
|
Reception
|
normal
|
ISR
|
|
| #1417 |
02 Apr 01:04 Calibration of ATRs and TRs
|
A. Coronado
|
Antennas
|
normal
|
ISR
|
|
| #1418 |
02 Abr 04:40 Fall Tx1 by variation Line AC
|
P. Condor
|
Transmitters
|
normal
|
ISR
|
|
| #1419 |
02 Apr 05:00 ChA is recovered after fall of TX1
|
P. Condor
|
Reception
|
normal
|
ISR
|
|
| #1420 |
02 Apr 05:18 Calibration of ATRs and TRs
|
A. Coronado
|
Antennas
|
normal
|
ISR
|
|
| #1421 |
02 Apr 00:00 Figures of RTI, noise, spectras, etc
|
P. Condor
|
Process/Analysis Program
|
normal
|
ISR
|
|
| #1422 |
02 Abr 09:03 Calibration of ATRs and TRs
|
J. Ortiz
|
Antennas
|
normal
|
ISR
|
|
| #1423 |
02 Apr 08:25 Signal Chain stop
|
J. Ortiz
|
Process/Analysis Program
|
normal
|
ISR
|
|
| #1424 |
02 Apr 10:01 Signal Chain restart
|
J. Ortiz
|
Process/Analysis Program
|
normal
|
ISR
|
|
| #1425 |
02 Apr 12:01 Drop in Tx line
|
J. Ortiz
|
Transmitters
|
normal
|
ISR
|
|
| #1426 |
02 Apr 12:22 Calibration of ATRs and TRs
|
J. Ortiz
|
Antennas
|
normal
|
ISR
|
|
| #1427 |
02 Apr 13:34 Posible Solar Flare Interference
|
J. Ortiz
|
Physical Phenomena
|
normal
|
ISR
|
|
| #1428 |
02 Apr 14:40 Power adjustment in Tx1 due to AC Line variation
|
J. Ortiz
|
Transmitters
|
normal
|
ISR
|
|
| #1429 |
02 Apr 15:20 Signal Chain Restart
|
J. Ortiz
|
Process/Analysis Program
|
normal
|
ISR
|
|
| #1430 |
02 Apr 15:22 Calibration of ATRs and TRs
|
J. Ortiz
|
Antennas
|
normal
|
ISR
|
|
| #1431 |
02 Apr 18:35 Calibration of ATRs and TRs
|
J. Ortiz
|
Antennas
|
normal
|
ISR
|
|
| #1432 |
02 Apr 19:50 Transmitter Power measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1433 |
03 Apr 00:12 Transmitter Power measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1434 |
02 Apr 21:55 Calibration of ATRs and TRs
|
jverastegui
|
Antennas
|
normal
|
ISR
|
|
| #1435 |
03 Apr 00:00 Figures of RTI, noise, spectras, etc
|
jverastegui
|
Process/Analysis Program
|
normal
|
ISR
|
|
| #1436 |
03 Apr 01:45 Calibration of ATRs and TRs
|
jverastegui
|
Antennas
|
normal
|
ISR
|
|
| #1437 |
03 Apr 07:10 Transmitter Power measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1439 |
03 Apr 07:05 Crowbar by voltage drop
|
jverastegui
|
Transmitters
|
normal
|
ISR
|
|
| #1440 |
02 Abr 07:50 Lose synchronization in the Faraday aquisition system
|
|
Acquisition system
|
normal
|
ISR
|
|
| #1441 |
03 Apr 08:41 Calibration of ATRs and TRs
|
dsuarez
|
General
|
normal
|
ISR
|
|
| #1442 |
03 Apr 11:53 Calibration of ATRs and TRs
|
dsuarez
|
General
|
normal
|
ISR
|
|
| #1443 |
03 Apr 15:09 Calibration of ATRs and TRs
|
dsuarez
|
General
|
normal
|
ISR
|
|
| #1444 |
03 Apr 17:03 IDL program is hanged
|
dsuarez
|
Process/Analysis Program
|
normal
|
ISR
|
|
| #1445 |
03 Apr 18:10 Calibration of ATRs and TRs
|
dsuarez
|
General
|
normal
|
ISR
|
|
| #1446 |
03 Apr 20:00 Transmitter Power mesearuments
|
|
Transmitters
|
normal
|
ISR
|
|
| #1447 |
03 Apr 22:20 spark gaps adjustment
|
I. Manay
|
Spark Gaps
|
normal
|
ISR
|
|
| #1448 |
04 Apr 23:50 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1449 |
03 Apr 23:30 Re-starting acquisition of the Faraday experiment
|
I. Manay
|
Acquisition system
|
normal
|
ISR
|
|
| #1450 |
03 Apr 23:57 Re-starting acquisition of the Faraday experiment
|
I. Manay
|
Acquisition system
|
normal
|
ISR
|
|
| #1451 |
03 Apr 23:59 plots of experiments
|
I. Manay
|
Plot
|
normal
|
ISR
|
|
| #1452 |
04 Apr 02:40 spark gaps adjustment
|
J. Lavado
|
Spark Gaps
|
normal
|
ISR
|
|
| #1453 |
04 Apr 06:55 spark gaps adjustment
|
J. Lavado
|
Spark Gaps
|
normal
|
ISR
|
|
| #1454 |
04 Apr 07:05 Transmitter Power Measurements
|
|
Transmitters
|
normal
|
ISR
|
|
| #1455 |
04 Apr 07:49 Re-starting acquisition of the Faraday experiment
|
I. Manay
|
Acquisition system
|
normal
|
ISR
|
|
| #1457 |
04 Abr 12:35 End of campaign
|
P. Condor
|
End experiment
|
normal
|
ISR
|
|
| #1458 |
09 Apr 13:30 Experiment data to PUMA
|
I. Manay
|
General
|
normal
|
ISR
|
|